two painters can paint a room in one work of six hours.
1.1 How long will the work shift be if four painters complete the work at the same rate and 1.2 if the room needs to be painted in a work shift of 2,4 hours, how many painters are then needed if they work at same rate as painters in 1.1

Answer by ankor@dixie-net.com(15747)

Here's the way I understand this:
:
two painters can paint a room in one shift of six hours.
determine how many man-hrs required to paint the room
2*6 = 12 man-hrs to complete the job
:
1.1 How long will the work shift be if four painters complete the work at the same rate?
let h = hrs required when 4 painter paint
4h = 12
h = 12/4
h = 3 hrs
:
1.2 if the room needs to be painted in a work shift of 2.4 hours, how many painters are then needed if they work at same rate as painters in 1.1
let m = no. of men required to do the job in 2.4 hrs
2.4m = 12
m = 12/2.4
m = 5 men required to accomplish this in 2.4 hrs